# Career Modeling with Missing Data and Traces

### v2

@@ 2.2.2 Formal presentation
For each individual $i$, we assume that the trajectory $(X_{t}^{i})_{t}$ is a Markov chain with non-homogeneous transition kernel $P_{i,t}=(\gamma_{i,t}^{x,x^{\prime}})_{x,x^{\prime}\in\mathcal{X}}$. For each starting point $x$, we model the transition probabilities using a generalized linear model (GLM) for the multinomial distribution:
@@
with $W_{i,t}$ the aforementioned vector of covariates, and $(\beta^{x,x^{\prime}})_{x^{\prime}\in\mathcal{X}}$ vectors of parameters. Notice that we may have different parameters $\beta^{x,x^{\prime}}$ for different transitions $(x,x^{\prime})\in\mathcal{X}\times\mathcal{X}$, but that the parameters $\beta^{x,x^{\prime}}$ do not depend on time. Time is rather encoded here as a variable in the covariate vectors $W_{i,t}$. Notice also that, for a state space of size $K$, there are only $K\times(K-1)$ vectors of parameters $(\beta^{x,x^{\prime}})$ to model. In the specific case of career modeling, where individuals do not move at every step, it is reasonable to take the absence of change $x=x^{\prime}$ as the reference category.
@@
with $\widetilde{W}_{i,t}$ the vector of covariates, $(\eta^{x})_{x\in\mathcal{X}}$ vectors of parameters, and $g$ a link function. We stress that we use the notation $\widetilde{W}_{i,t}$ to communicate that when estimating emission probabilities, it is possible to include lagged values of $(Y_{i,t^{\prime}})_{t^{\prime}<t}$ inside $\widetilde{W}_{i,t}$, in addition to the standard covariates in ${W}_{i,t}$. Formally, $\widetilde{W}_{i,t}=[W_{i,t}^{T},f(Y_{i,1},...,Y_{i,t-1})]^{T}$, with $f$ any user-specified transformation. We discuss why this auto-correlation feature is essential for such applications in subsection 2.2.3.
@@
$\forall x\in\mathcal{X},\mathcal{L}(\beta^{x,.}|\textbf{X, W})=\prod_{i=1}^{N}\prod_{\begin{subarray}{c}t=1\\ X_{i,t}=x\end{subarray}}^{T-1}\prod_{x^{\prime}\in\mathcal{X}}\left(\gamma_{i,t}^{x,x^{\prime}}\right)^{1_{\{X_{i,t+1}=x^{\prime}\}}}$ (3)
@@
$\forall x\in\mathcal{X},\mathcal{L}(\eta^{x}|\textbf{X, Y, W})=\prod_{i=1}^{N}\prod_{\begin{subarray}{c}t=1\\ X_{i,t}=x\end{subarray}}^{T}\mathcal{H}(Y_{i,t}|\lambda_{i,t}^{x}),$ (4)
@@
where $\mathcal{H}$ is the likelihood function for the trace-emission GLM model, which will be practical for exploring the space of parameters with a custom sampler.
We purposefully introduced the Markov switching model in a general fashion, but it is now useful to provide details on one useful specificity of applications for career modeling. In many instances, certain values taken by $X_{i,t}\in\mathcal{X}$ will be tied to certain values or a certain support of $Y_{i,t}$, deterministically. For example: when studying traces in the public sector record, we know that if an individual is working in the private sector (denoted here as $X_{i,t}=\text{Private}$) then there cannot be any traces in the public sector record. Formally, $X_{i,t}=\text{Private}$ implies $Y_{i,t}=0$. This can be applied to many cases where individuals cannot hold multiple positions at the same time: journalists who have a full-time contract, managers in companies, politicians and their political party.
@@
There is no need to actively model the emission parameters $\eta^{x}$ in cases $x\in\mathcal{X}$ where we have such information (e.g. $\eta^{Private}$ for the civil servants), since, because of our external knowledge, the conditional distribution of $Y_{i,t}$ given $X_{i,t}=x$ is deterministic. This reduces the number of parameters to estimate and enables the seamless integration of expert knowledge. Additionally, when trying to reconstruct trajectories X from traces Y, this will provide an unusually strong conditioning compared to usual applications of HMMs. This is a reason why our procedure works well in practice. As a whole, this application specificity thus facilitates both inference of parameters and reconstruction of trajectories.
2.2.3 Motivation from a complexity point of view
@@
The model we suggest is purposefully more complex than a Hidden Markov Model (HMM). Indeed, several issues arise from simply using a HMM in our application. One is substantive: a key interest is to describe how career behavior varies across individuals, groups, and time; thus we naturally need to include covariates in the models. Another is more technical and concerns traces. A key component of the problem is that it is essential that the trace-emission process be well modeled: otherwise, errors here will be communicated to the rest of the procedure, which will result both in a poor reconstruction of X and in biased inference of the $(\beta^{x,x^{\prime}})$. We present for now the substantive reason for why it is essential to have covariates and lagged values of traces when modeling the trace-emission process. We defer examples based on synthetic data to section 3.2 as they require more notation.
@@
Two problems make the inclusion of covariates and lagged values essential in modeling the trace-emission process: heterogeneity and self-excitation. In all the examples cited above, we expect heterogeneity between individuals emitting traces. Some journalists and academics write more articles than others, some civil servants make more important decisions and are promoted more often than their peers, and some athletes participate in more competitions than their competitors. If we do not account for this, our inference will confound the trace frequency variety and the actual behavior of careers. Furthermore, in most of the examples we cited, it is expected that some traces will be tied or will lead to other traces in quick succession, a fact that can either be described as self-excitation or time heterogeneity. Politicians writing political posts will lead to more posting, civil servants joining a committee will lead to more announcements when the committee arrives at a position, and journalists writing a piece on some hot-button issue can be expected to write more in the following weeks or months. Not accounting for that will lead to overestimating the probability of finding traces at any given time, and thus to overestimating the number of transitions across states of trajectories X. For this reason, using a simple HMM is not sufficient, and there is a need for both covariates and transformations of lagged values of traces.
On the other hand, it might be argued that a more complex model is called for. We resist the urge to go for more complex models, because Markov switching models exhibit properties that greatly facilitate the inference and enable us to efficiently target the posterior in a Bayesian framework. In particular, their factorization structure still enables the usage of the Baum-Welch algorithm when reconstructing sequences X from traces Y. One can take any presentation of HMM reconstruction (e.g. Marin and Robert, 2014), simply subscript the transition and emission rates, and perform the reconstruction unchanged. In fact, this would even work for rates that are obtained with more complex models than a GLM, such as neural networks (Tran et al., 2016), as long as the independence structure described in Figure 1 holds. We do not consider such extensions here, and remain in the GLM realm as our goal is to test substantive hypotheses.
This last point connects to one final discussion of the goals of the analysis. If the goal were only to reconstruct the trajectories X with the highest accuracy at the individual level, other solutions from (Bayesian) machine learning would probably be better suited, using recent developments in deep learning for sequences and inverse problems. However, our primary goal here is to provide a method that enables the testing of substantive research hypotheses, reconstruction being only secondary and used in an MCMC sampler. We still report quantities related to reconstruction accuracy with cross-validation experiments, but our main interest throughout is to enable the unbiased and precise inference of substantively motivated parameters. Yet, more complex developments could be motivated for the trace emission part (since we consider emission parameters $\eta^{x}$ to be nuisance), and for the career trajectory part (notably using varying-length Markov models, see e.g. Le Picard and Rossi, 2023).
2.2.4 MCMC inference sampler
@@
We propose a tailored estimation scheme by MCMC to target the posterior, consisting of one Gibbs step for generating the missing values of X, and Metropolis-within-Gibbs steps for exploring the parameter spaces. This contrasts with a recent work from Michelot (2022) which proposes a faster (but non-exact) estimation using the Laplace approximation of the posterior. We give here a general presentation on our algorithm, but note that our implementation (available at https://github.com/theovoldoire/Hidden_Careers_Public) is more problem-specific: in our implementation, we only consider a state space of dimension 2, and perform several simplifications similar to those discussed in section 2.2.2 and introduced in section 3. The extension to the general case would be straightforward.
@@
Gibbs step for unobserved values of $X$. The Gibbs step consists of simulating missing values of X given observed values of X, traces Y, and parameters $(\beta^{x,x^{\prime}})_{x,x^{\prime}\in\mathcal{X}}$, $(\eta^{x})_{x\in\mathcal{X}}$, and consists of two stages. In the first stage, we compute the transition parameters $\gamma_{i,t}^{x,x^{\prime}}$ for all individuals $i$, time $t$, and states $(x,x^{\prime})\in\mathcal{X}\times\mathcal{X}$; and the conditional probabilities $p(Y_{i,t}|X_{i,t}=x,\widetilde{W}_{i,t},\eta^{x})$ provided by the expected value $\lambda_{i,t}=\mathbb{E}[Y_{i,t}|X_{i,t}=x,\widetilde{W}_{i,t},\eta^{x}]=g(\widetilde{W}_{i,t}^{T}\eta^{x})$, for all $i,t$ and states $x$. In the second stage, we use the Baum-Welch algorithm. Our implementation is based on the R package HMM (Himmelmann, 2022), for which we simply added a time subscript on parameters; note that the Baum-Welch algorithm allows for transition and emission parameters which are not constant across time and units.
Metropolis-within-Gibbs step for parameters. We take inspiration from Marin and Robert (2014) and apply the same method successively to different parameters of interest. Iterating on $x\in\mathcal{X}$, we update separately each vector of parameters, first $(\beta^{x,x_{1}},...,\beta^{x,x_{K}})$ jointly, and then $\eta^{x}$. For each update of an arbitrary parameter vector $\mu$, we proceed as follows: compute the MLE $\hat{\mu}$ for this given parameter using either equation 3 or 4 and compute the associated covariance matrix $\hat{\Sigma}$. Second, run for $M_{\text{metropolis}}$ iterations a Metropolis-Hastings random walk with the proposition $\tilde{\mu}\sim\mathcal{N}(\mu,\tau^{2}\hat{\Sigma})$, with $\tau$ a hyperparameter that we set to 1 after experimentation. After $M_{\text{metropolis}}$ iterations, keep the final value of parameter $\mu$ as the new value. In our applications, visual checks indicated that it was sufficient to take $M_{\text{metropolis}}=30$.

### v3

@@ 3.2 Formal presentation
For each individual $i$, we assume that the trajectory $(X_{t}^{i})_{t}$ is a Markov chain over with non-homogeneous transition kernel $P_{i,t}=(\gamma_{i,t}^{x,x^{\prime}})_{x,x^{\prime}\in\mathcal{X}}$. For each starting point $x\in\mathcal{X}$, we model the transition probabilities using a generalized linear model (GLM) with multinomial distribution:
@@
with $W_{i,t}$ the aforementioned vector of covariates, and $(\beta^{x,x^{\prime}})_{x^{\prime}\in\mathcal{X}}$ vectors of parameters. In this regression, we take $x\to x$ as the reference category and so set $\forall x\in\mathcal{X},\beta^{x,x}=0$. Notice that we may have different parameters $\beta^{x,x^{\prime}}$ for different transitions $(x,x^{\prime})\in\mathcal{X}\times\mathcal{X}$, but that the parameters $\beta^{x,x^{\prime}}$ do not depend on time. Time is rather encoded here as a variable in the covariate vectors $W_{i,t}$. It is straightforward to extend this to the multinomial case when the state space $\mathcal{X}$ is of cardinal greater than $2$. Notice that, for a state space of size $K$, there are only $K\times(K-1)$ vectors of parameters $(\beta^{x,x^{\prime}})$ to model. In the specific case of career modeling, where individuals do not move at every step, it is reasonable to take the absence of change $x=x^{\prime}$ as the reference category.
@@
with $\widetilde{W}_{i,t}$ the vector of covariates, $(\eta^{x})_{x\in\mathcal{X}}$ vectors of parameters, and $g$ a link function. We stress that we use the notation $\widetilde{W}_{i,t}$ to communicate that when estimating emission probabilities, it is possible to include lagged values of $(Y_{i,t^{\prime}})_{t^{\prime}<t}$ inside $\widetilde{W}_{i,t}$, in addition to the standard covariates in ${W}_{i,t}$. Formally, $\widetilde{W}_{i,t}=[W_{i,t}^{T},f(Y_{i,1},...,Y_{i,t-1})]^{T}$, with $f$ any user-specified transformation. We discuss why this auto-correlation feature is essential for our application in subsection 3.3.
[…]
@@
$\forall x\in\mathcal{X},\mathcal{L}(\beta^{x,.}|\textbf{X, W})=\prod_{i=1}^{N}\prod_{\begin{subarray}{c}t=1\\ \text{s.t. }X_{i,t}=x\end{subarray}}^{T-1}\prod_{x^{\prime}\in\mathcal{X}}\left(\gamma_{i,t}^{x,x^{\prime}}\right)^{1_{\{X_{i,t+1}=x^{\prime}\}}}$ (3)
@@
$\forall x\in\mathcal{X},\mathcal{L}(\eta^{x}|\textbf{X, Y, W})=\prod_{i=1}^{N}\prod_{\begin{subarray}{c}t=1\\ \text{s.t. }X_{i,t}=x\end{subarray}}^{T}\mathcal{H}(Y_{i,t}|\lambda_{i,t}^{x}),$ (4)
@@
where $\mathcal{H}$ is the likelihood function for the trace-emission GLM model, which will be practical for exploring the space of parameters with a custom sampler. In both equations, the second product is only over time indices $t$ such that $X_{i,t}=x$.
@@
We purposefully introduced the Markov switching model in a general fashion, but our career modeling application presents a useful specificity: certain values taken by $X_{i,t}\in\mathcal{X}$ are tied to certain values or a certain support of $Y_{i,t}$, deterministically. Indeed, we know that if an individual is working in the private sector (denoted here as $X_{i,t}=\text{Private}$) then there cannot be any traces in the public sector record. Formally, $X_{i,t}=\text{Private}$ implies $Y_{i,t}=0$. There is no need to actively model the emission parameters $\eta^{x}$ in cases $x\in\mathcal{X}$ where we have such information, since, because of our external knowledge, the conditional distribution of $Y_{i,t}$ given $X_{i,t}=x$ is deterministic. This reduces the number of parameters to estimate and enables the seamless integration of expert knowledge. Additionally, when trying to reconstruct trajectories X from traces Y, this will provide an unusually strong conditioning compared to usual applications of HMMs. This is a reason why our procedure works well in practice. As a whole, this application specificity thus facilitates both inference of parameters and reconstruction of trajectories.
3.3 Motivation from a complexity point of view
@@
The model we suggest is purposefully more complex than a Hidden Markov Model (HMM). Indeed, several issues arise from simply using a HMM in our application. One is substantive: a key interest is to describe how career behavior varies across individuals, groups, and time; thus we naturally need to include covariates in the models. Another is more technical and concerns traces. A key component of the problem is that it is essential that the trace-emission process be well modeled: otherwise, errors here will be communicated to the rest of the procedure, which will result both in a poor reconstruction of X and in biased inference of the $(\beta^{x,x^{\prime}})$.
@@
Two problems make the inclusion of covariates and lagged values essential in modeling the trace-emission process: heterogeneity and self-excitation. We expect heterogeneity between individuals emitting traces, for instance because some civil servants make more important decisions and are promoted more often than their peers. If we do not account for this, our inference will confound the trace frequency variety and the actual behaviour of careers. Furthermore, it is expected that some traces will be tied or will lead to other traces in quick succession, a fact that can either be described as self-excitation or time heterogeneity: for example, civil servants joining a committee will lead to more announcements when the committee arrives at a position. Not accounting for that will lead to overestimating the probability of finding traces at any given time, and thus to overestimating the number of transitions across states of trajectories X. For this reason, using a simple HMM is not sufficient, and there is a need for both covariates and transformations of lagged values of traces. We show on synthetic data in Section 4 that a misspecified HMM does lead to biased inference. Our results on real data in Section 5 confirm that this choice is appropriate: when we incorporate lagged values and gender, cohort, and organization covariates, we obtain better model fit (lower reconstruction loss) and find that many of the corresponding parameters are significantly different from 0.
3.4 MCMC inference sampler
Our framework relies on a large number of latent variables, making it natural to perform inference via data augmentation. Uncertainty quantification is essential: for instance, the career paths of certain individuals will remain highly uncertain, and we need this to translate into uncertainty about the corresponding model parameters. More generally, we wish to allow for uncertainty to flow between the model components, and between the two data sources. We thus perform inference in a Bayesian setting with vague priors. We describe here our algorithm to sample from the joint posterior of all model parameters and latent variables.
@@ 3.4 MCMC inference sampler
We propose a tailored estimation scheme by MCMC to target the posterior, consisting of one Gibbs step for generating the missing values of X, and Metropolis-within-Gibbs steps for exploring the parameter spaces. This contrasts with a recent work from michelot_hmmtmb_2023 which proposes a faster (but non-exact) estimation using the Laplace approximation of the posterior. We give here a general presentation on our algorithm, but note that our implementation (available at https://github.com/robinryder/career-paths-public) is problem-specific: in our implementation, we only consider a state space of dimension 2, and perform several simplifications similar to those discussed in section 3.2 and introduced in section 5. The extension to the general case would be straightforward.
@@
Gibbs step for unobserved values of $X$. The Gibbs step consists of simulating missing values of X given observed values of X, traces Y, and parameters $(\beta^{x,x^{\prime}})_{x,x^{\prime}\in\mathcal{X}}$, $(\eta^{x})_{x\in\mathcal{X}}$, and consists of two stages. In the first stage, we compute the transition parameters $\gamma_{i,t}^{x,x^{\prime}}$ for all individuals $i$, time $t$, and states $(x,x^{\prime})\in\mathcal{X}\times\mathcal{X}$; and the conditional probabilities $p(Y_{i,t}|X_{i,t}=x,\widetilde{W}_{i,t},\eta^{x})$ provided by the expected value $\lambda_{i,t}=\mathbb{E}[Y_{i,t}|X_{i,t}=x,\widetilde{W}_{i,t},\eta^{x}]=g(\widetilde{W}_{i,t}^{T}\eta^{x})$, for all $i,t$ and states $x$. In the second stage, we use the Baum-Welch algorithm. Our implementation is based on the R package HMM (HMM), for which we simply added a time subscript on parameters; note that the Baum-Welch algorithm allows for transition and emission parameters which are not constant across time and units.
[…]
In practice, the first step (data augmentation) takes up the vast majority of the computational time.
4 Empirical demonstrations on synthetic data
We present empirical demonstrations of multiple claims made in Section 2, taking advantage of the notation and intuitions that arise from this specific case. We first show that the conjunction of detailed data and trace data leads to more accurate statistical inference: compared to using only detailed data, our credible intervals are substantially narrower. We then demonstrate a limitation of the (simpler) standard HMM framework: even if we allow an HMM with heterogeneity in the trace-emitting process, a misspecified model leads to errors which propagate and to inferential biases; our approach can immunize against this problem.
4.1 Gains in statistical accuracy
We first demonstrate that information is gained by considering the two data sources, rather than only the detailed data. We simulate trajectories using the following form for parameters, with $g^{-1}$ the inverse logit transform:
$\forall i,t,\ \gamma^{0\rightarrow 1}_{i,t}=g^{-1}(\alpha_{0}+\beta_{0}X_{i})\qquad\ \gamma^{1\rightarrow 0}_{i,t}=g^{-1}(\alpha_{1})\qquad\lambda_{i,t}=g^{-1}(\alpha_{\lambda}),$ (5)
with $X_{i}\sim\mathcal{U}([0,1])$ an individual observed covariable, that could correspond in our example to the cohort of individual $i$. Our main parameter of interest is the parameter $\beta_{0}$ which measures the impact of the cohort covariable on the probability for individuals to leave the public sector.
We simulated synthetic data using the parameter values $\alpha_{0}=-4$, $\beta_{0}=1$, $\alpha_{1}=-4$, $\alpha_{\lambda}=-1.5$. These coefficients are similar to estimates provided for the simplest models in Section 5.2, so this simulation experiment can be deemed realistic. Each trajectory is initiated at state 0, and is of length 60. We simulated three datasets: (1) $N=200$ individuals, all observed in full in the detailed source; (2) $N=500$ individuals, all observed in full in the detailed source; (3) $N=1000$ individuals, of which only $n_{obs}=200$ are observed in the detailed source but all of which are present in the trace data. Note that in cases 1 and 2, all paths are fully observed: the inference procedure is straightforward, and we do not expect any gain from the trace data. In case 3, we get the same number of fully observed individuals as in case 1, but we also get trace data for more individuals. This synthetic data experiment allows us to quantify the amount of information we gain from the trace data.
Table 1: Posterior 90% credible intervals for the synthetic data experiment of Section 4.1.
Case 1 Case 2 Case 3 True value
NN 200 500 1000
$n_{\text{obs}}$ 200 500 200
$\alpha_{0}$ [-4.22,-3.51] [-4.26,-3.81] [-4.19,-3.8] -4
$\beta_{0}$ [-1.04,1.25] [0.17,1.64] [0.72,1.96] 1
$\alpha_{1}$ [-4.91,-4.04] [-4.11,-3.69] [-4.36,-3.81] -4
$\alpha_{\lambda}$ [-1.54,-1.41] [-1.52,-1.44] [-1.5,-1.44] -1.5
We use the MCMC algorithm described in Section 3.4 to target the joint posterior distribution of $(\alpha_{0},\beta_{0},\alpha_{1},\alpha_{\lambda})$. We report on the concentration of the posterior of coefficients, by reporting on 90% credibility intervals in Table 1, from which different conclusions can be drawn. First, compare case 3 to case 1, and recall that these two cases have the same number of detailed observations but case 3 also includes traces: the inference in case 3 is better across the board; all 90% posterior intervals contain the true value, but the intervals in case 3 are narrower than those in case 1, by a factor between 1.6 and 2.2 depending on the parameter of interest. Second, cases 2 and 3 can be said to have relatively similar performances, with comparable posteriors and uncertainty. However, the gain here is that while case 2 relies on observing 500 individuals, in case 3 only 200 individuals had to be observed in the detailed data source, and imprecise knowledge using traces compensated precise knowledge on a population more than twice larger. In practice, we expect there to be a constraint on how large $n_{obs}$ can be – either because of data availability, or because it is costly to collect detailed data on many individuals. This simulation study shows that with trace data on $N=1000$ individuals and detailed data on only $n_{obs}=200$ individuals, we get similar statistical accuracy than if we had detailed data on $N=n_{obs}=500$ individuals, a gain of a factor of 2.5 in the cost of collecting the detailed data.
4.2 Issues with standard HMM misspecification
We now consider a situation where a standard HMM is misspecified. We show that inference with a standard HMM leads to the misspecification error propagating to other inferred parameters. We simulate trajectories using the following form for parameters
$\forall i,t,\ \gamma^{0\rightarrow 1}_{i,t}=g^{-1}\left(\alpha_{0}+\beta_{0}B_{i,t}\right),\ \gamma^{1\rightarrow 0}_{i,t}=g^{-1}(\alpha_{1}),\ \lambda_{i,t}=g^{-1}(\alpha_{\lambda}+\beta_{\lambda}A_{i,t}(0.8)),$ (6)
where $B_{i,t}=\frac{t}{60}$ is a time-varying covariate, which here could correspond to period, and $A_{i,t}(0.8)$ the autocorrelation covariate as defined earlier. This is a realistic scenario: as mentioned above, we don’t expect traces to be emitted independently. Instead, we expect some positive autocorrelation: when an individual emits a trace, we expect them to have a higher probability of emitting further traces at the next time steps. We simulate synthetic data with the values $\alpha_{0}=-4$, $\beta_{0}=1$, $\alpha_{1}=-4$, $\alpha_{\lambda}=-2$, $\beta_{\lambda}=2.5$. We simulate 2000 sequences of length 60; we observe detailed data for 30% of the individuals, and trace data for all. We implement two models: one well-specified (ours), and one misspecified (standard HMM), where covariate $A_{i,t}$ is not included, i.e. $\beta_{\lambda}=0$.
Table 2: Posterior 90% credible intervals for the synthetic data experiment of Section 4.2
Well-specified Misspecified True value
$\alpha_{0}$ [-4.28,-3.99] [-3.72,-3.49] -4
$\beta_{0}$ [0.86,2.03] [-0.32,0.66] 1
$\alpha_{1}$ [-4.37,-3.95] [-4.09,-3.78] -4
$\alpha_{\lambda}$ [-2.04,-1.96] [-1.49,-1.45] -2
$\beta_{\lambda}$ [2.39,2.61] 2.5
We sample from the joint posterior distribution of all model parameters. Coefficients are reported in Table 2. In the well-specified case, all the output credibility intervals correctly contain the true parameter values. On the other hand, in the misspecified HMM case, multiple issues appear. Not only is there no estimate for the parameter $\beta_{\lambda}$, but the misspecification propagates: the posterior intervals for $\beta_{0}$ and $\alpha_{\lambda}$ are far from including the true values. Such problems will arise for time-varying covariates when an autocorrelation behavior is not taken into account in the modeling phase. The same point is also shown on real data in Section 5.2.2.